OpenStudy (anonymous):

Can anyone explain this to me? 4/7 = x+3/21 Thanks :)

OpenStudy (anonymous):

(x+3)/21?

OpenStudy (anonymous):

or (3/21)?

OpenStudy (anonymous):

replace 3/21 to be 1/7 so x = 3/7

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

its not one of the answers :/ I am usually pretty good at math but I don't like the way they phased this question. the answer choices are A. x=9 B. x=91 C. x=33 (3/4) D. x=15

OpenStudy (anonymous):

(x-3)/21 @DemolisionWolf

OpenStudy (anonymous):